A coffee shop blends a coffee that sells for $12 per pound with a coffee that sells for $9 per pound to produce a mixture that sells for $10 per pound. How much of each should be used to yield 21 pounds of the new mixture?
Answer by lwsshak3(11628)

A coffee shop blends a coffee that sells for $12 per pound with a coffee that sells for $9 per pound to produce a mixture that sells for $10 per pound. How much of each should be used to yield 21 pounds of the new mixture?
**
let x=lbs of $12/lb coffee to be mixed
21-x=lbs of $9/lb coffee to be mixed
..
12x+9(21-x)=10(21)
12x+189-9x+=210
3x=21
x=7
21-x=14
lbs of $12/lb coffee to be mixed=7
lbs of $9/lb coffee to be mixed=14